# Frobenius pull-back of parabolic bundles and dormant opers

### v1

@@ 6.1. Maximally $F^{(N)}$-destabilized bundles
    The inequalities $\sum\limits_{i=1}^{r}(s_{1}^{1}(a_{i}^{[n]})-s_{1}^{1}(a_{i}^{[1]}))<\frac{n(2g-2+r)}{2}\leq\frac{p}{n}$ hold.
@@
Assertion (i) can be proved by a routine argument using $02g-2+r>$ 0 (so we omit the detail of the proof).
[…]
@@
Now, suppose that $\mathcal{F}^{j}$ (for some $j=1,\cdots,n-1$) of $\mathcal{F}$ is closed under $\nabla^{F}$. If $\nabla^{j}$ denotes the log connection on $\mathcal{F}^{j}$ obtained by restricting $\nabla^{F}$, then we have $F^{*}_{X/S}\mathcal{S}ol(\nabla^{j})=F^{*}_{X/S}\mathcal{S}ol(\nabla)\cap\mathcal{F}^{j}$ (where $F_{X/S}:=F^{(1)}_{X/S}$), and both sides of this equality have degree divided by $p$. Let us observe the following equalities:
@@
where the inequality “$<$” follows from the first inequality in the condition (c) and the last inequality $\leq$” follows from the second one in (c). On the other hand, we have
@@ 6.2. Correspondence with dormant $\mathrm{GL}_{n}^{(N)}$-opers
Assertion (i) follows from Proposition 5.6, (ii).
We shall consider assertion (ii). The assignment $\mathscr{E}\mapsto\mathscr{E}^{F,\heartsuit}$ is functorial because of its construction, so it defines a functor
@@
We shall construct its inverse. Let $(\mathscr{F}^{\heartsuit},\eta)$ (where :=(,,$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$) be an object classified by $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}^{F}}$. Apply Lemma 6.5 described below to the restriction of $(\mathcal{F},\nabla)$ to the formal neighborhood of each $\sigma_{i}$. Then, there exists a quasi-parabolic structure $\vec{\mathpzc{f}}$ on $\mathcal{F}$, for which $(\mathcal{F},\nabla,\vec{\mathpzc{f}},\vec{a})$ specifies a parabolic $p^{N}$-flat bundle. It follows from Theorem 4.5 that the horizontal parabolic bundle $\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\vec{\mathpzc{f}}^{\nabla},\vec{a}/p^{N})$ associated to this collection is maximally $F^{(N)}$-destabilized. Moreover, by the assumptions on $a_{i}^{[j]}$’s imposed in (ii), we see that $\mathscr{F}^{\nabla}$ belongs to $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}(S)$ (cf. Proposition 6.2, (ii)). The formation of the assignment $(\mathscr{F}^{\heartsuit},\eta)\mapsto\mathscr{F}^{\nabla}$ is functorial with respect to $S$, and hence, determines a functor
@@
Let $(\mathcal{F},\nabla)$ be a rank-$n$ $p^{N}$-flat bundle on $\mathscr{U}$ whose exponent is given by $[a_{\circ}^{[1]},\cdots,a_{\circ}^{[n]}]$ for some element $a_{\circ}:=(a_{\circ}^{[1]},\cdots,a_{\circ}^{[n]})$ of $\Xi_{n,N}^{<}$. Then, there exists a unique quasi-parabolic structure $\mathpzc{f}_{\circ}$ on $\mathcal{F}$ for which $(\mathcal{F},\nabla,\mathpzc{f}_{\circ})$ forms a parabolic $p^{N}$-flat bundles.
@@
Denote by $d^{[j]}$ ($j=1,\cdots,n$) the image of $a_{\circ}^{[j]}$ via the quotient $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$. To prove this assertion, we may assume, without loss of generality, that there exists an isomorphism of $\mathcal{D}_{\circ}^{(N-1)}$-modules $\xi:(\mathcal{F},\nabla)\xrightarrow{\sim}\bigoplus_{j=1}^{n}\mathscr{O}_{d^{[j]},\flat}$. By considering the eigenspaces of the monodromy operator of $\bigoplus_{j=1}^{n}\nabla_{d^{[j]}}$ at $\sigma_{\circ}$, it is verified that there exists a unique quasi-parabolic structure $\mathpzc{f}^{\prime}_{\circ}$ on $\mathcal{O}_{U}^{\oplus n}$ for which the collection $(\mathcal{O}_{U}^{\oplus n},\bigoplus_{j=1}^{n}\nabla_{d^{[j]}},\mathpzc{f}^{\prime}_{\circ})$ forms a parabolic $p^{N}$-flat bundle. The quasi-parabolic structure $\mathpzc{f}_{\circ}$ on $\mathcal{F}$ corresponding to $\mathpzc{f}^{\prime}_{\circ}$ via $\xi$ provides a required one.
Moreover, since $a_{\circ}^{[1]}<\cdots<a_{\circ}^{[n]}$, it follows from the definition of $\mathpzc{f}^{\prime}_{\circ}$ and  [Wak4, Proposition 4.3.4, (ii)] that $\mathpzc{f}^{\prime}_{\circ}$ is invariant under transposing by any automorphism of $\bigoplus_{i=1}^{n}\mathscr{O}_{d_{\circ}^{[j]},\flat}$. Hence, $\mathpzc{f}_{\circ}$ does not depend on the choice of $\xi$. This proves the uniqueness assertion, as desired. ∎
@@
Let us keep the above notation and assumptions, and suppose that $n<p$ and $\sum\limits_{j=1}^{n}a_{i}^{[j]}<p^{N}$ for every $i=1,\cdots,r$. Then, the following assertions hold.
@@
Let us take a line bundle $\mathcal{L}$ on $X^{(1)}$ of even degree and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in(\Xi_{2,1}^{<})^{\times r}$, where $a_{i}:=(a_{i}^{[1]},a_{i}^{[2]})$ ($i=1,\cdots,r$). We write $\omega:=(2,\vec{1}_{\times 2},\vec{a}/p)$, and suppose that $r+\sum\limits_{i=1}^{r}(a_{i}^{[1]}+a_{i}^{[2]})$ is even and $\sum\limits_{i=1}^{r}(a_{i}^{[2]}-a_{i}^{[1]})<2g-2+r\leq\frac{p}{2}$. Then, the cardinality $\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))$ of the set of isomorphism classes of objects in $\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k)$ satisfies the following inequality:
@@
where, for an integer $b$ with $10\leq b\leq p-$ 1, we set $\tau(b):=\frac{b-1}{2}$ (resp., $\tau(b):=\frac{p-1-b}{2}$) if $b$ is odd (resp., even). On the other hand, the following equality holds:

### v2

@@ 6.1. Maximally $F^{(N)}$-destabilized bundles
    The inequalities $\sum\limits_{i=1}^{r}(s_{1}^{1}(a_{i}^{[n]})-s_{1}^{1}(a_{i}^{[1]}))<\frac{n(2g-2+r)}{2}<\frac{p}{n}$ hold.
@@
Assertion (i) can be proved by a routine argument using 2g2+r>02g-2+r>0 (so we omit the detail of the proof).
In the following, we prove assertion (ii). Suppose that there exists a proper parabolic subbundle $\mathscr{G}:=(\mathcal{G},\vec{\mathfrak{g}},\vec{a}/p^{N})$ of $\mathscr{E}$ satisfying $\mathrm{par}\text{-}\mu(\mathscr{G})\geq\mathrm{par}\text{-}\mu(\mathscr{E})$. By Proposition 4.6, the parabolic Frobenius pull-back $\mathscr{G}^{F}:=(\mathcal{G}^{F},\nabla^{F}_{\mathcal{G}},\vec{\mathfrak{g}}^{F},\vec{a})$ specifies a parabolic $p^{N}$-flat subbundle of $\mathscr{E}^{F}$. Denote by $m\left(<n\right)$ the rank of $\mathcal{G}^{F}$, and define a decreasing filtration $\{\mathcal{G}^{F,j}\}_{j=0}^{n}$ on $\mathcal{G}^{F}$ by putting $\mathcal{G}^{F,j}:=\iota^{-1}(\mathcal{F}^{j})$, where $\{\mathcal{F}^{j}\}_{j=0}^{n}$ denotes the destabilized filtration of $\mathscr{E}$ and $\iota$ denotes the natural inclusion $\mathcal{G}^{F}\hookrightarrow\mathcal{F}\left(=\mathcal{F}^{0}=\mathcal{E}^{F}\right)$. Then, one obtains a subset $B:=\left\{j\,|\,\mathcal{G}^{F,j}\neq\mathcal{G}^{F,j+1}\right\}$ of $\{0,\cdots,n-1\}$. By Lemma 6.3 below together with an argument similar to the proof of Proposition 5.6 (by using the assumption (b)), we see that the log connection on $\mathcal{F}$ induced from $\nabla^{F}$ gives, in a usual manner, an injection $\mathcal{F}^{j}/\mathcal{F}^{j+1}\hookrightarrow\Omega\otimes(\mathcal{F}^{j-1}/\mathcal{F}^{j})$ (i.e., the $j$-th Kodaira-Spencer map). By the condition (a), this morphism is verified to be an isomorphism; it moreover restricts to an injection $\mathcal{G}^{F,j}/\mathcal{G}^{F,j+1}\hookrightarrow\Omega\otimes(\mathcal{G}^{F,j-1}/\mathcal{G}^{F,j})$. It follows that $j\in B\cap\{1,\cdots,n-1\}$ implies $j-1\in B$, and hence that $B=\{0,\cdots,m-1\}$. Hence, the composite $\mathcal{G}^{F}\xrightarrow{\mathrm{inclusion}}\mathcal{F}\twoheadrightarrow\mathcal{F}/\mathcal{F}^{m}$ turns out to be injective. This implies
@@
[…]
@@
where the inequality “<<” follows from the first inequality in the condition (c) and the last inequality $\leq$” follows from the second one in (c). On the other hand, we have
@@ 6.2. Correspondence with dormant $\mathrm{GL}_{n}^{(N)}$-opers
First, we shall consider assertion (i). By Proposition 6.2, (ii), the parabolic Frobenius pull-back of $\mathscr{E}$ by $F_{X/S}^{(N-1)}$ is stable. Hence, the assertion follows from Proposition 5.6, (ii).
Next, we shall consider assertion (ii). The assignment $\mathscr{E}\mapsto\mathscr{E}^{F,\heartsuit}$ is functorial because of its construction, so it defines a functor
@@
We shall construct its inverse. Let $(\mathscr{F}^{\heartsuit},\eta)$ (where $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$) be an object classified by $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}^{F}}$. Apply Lemma 6.5 described below to the restriction of $(\mathcal{F},\nabla)$ to the formal neighborhood of each $\sigma_{i}$. Then, there exists a quasi-parabolic structure $\vec{\mathfrak{f}}$ on $\mathcal{F}$, for which $(\mathcal{F},\nabla,\vec{\mathfrak{f}},\vec{a})$ specifies a parabolic $p^{N}$-flat bundle. It follows from Theorem 4.5 that the horizontal parabolic bundle $\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\vec{\mathfrak{f}}^{\nabla},\vec{a}/p^{N})$ associated to this collection is maximally $F^{(N)}$-destabilized. Moreover, by the assumptions on $a_{i}^{[j]}$’s imposed in (ii), we see that $\mathscr{F}^{\nabla}$ belongs to $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}(S)$ (cf. Proposition 6.2, (ii)). The formation of the assignment $(\mathscr{F}^{\heartsuit},\eta)\mapsto\mathscr{F}^{\nabla}$ is functorial with respect to $S$, and hence, determines a functor
@@
Let $(\mathcal{F},\nabla)$ be a rank-$n$ $p^{N}$-flat bundle on $\mathscr{U}$ whose exponent is given by $a_{\circ}^{[1]},\cdots,a_{\circ}^{[n]}]$ for some element $a_{\circ}:=(a_{\circ}^{[1]},\cdots,a_{\circ}^{[n]})$ of $\Xi_{n,N}^{<}$. Then, there exists a unique quasi-parabolic structure $\mathfrak{f}_{\circ}$ on $\mathcal{F}$ for which $(\mathcal{F},\nabla,\mathfrak{f}_{\circ})$ forms a parabolic $p^{N}$-flat bundles.
@@
Denote by $d^{[j]}$ ($j=1,\cdots,n$) the image of $a_{\circ}^{[j]}$ via the quotient $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$. To prove this assertion, we may assume, without loss of generality, that there exists an isomorphism of $\mathcal{D}_{\circ}^{(N-1)}$-modules $\xi:(\mathcal{F},\nabla)\xrightarrow{\sim}\bigoplus_{j=1}^{n}\mathscr{O}_{d^{[j]},\flat}$. By considering the eigenspaces of the monodromy operator of $\bigoplus_{j=1}^{n}\nabla_{d^{[j]}}$ at $\sigma_{\circ}$, it is verified that there exists a unique quasi-parabolic structure $\mathfrak{f}^{\prime}_{\circ}$ on $\mathcal{O}_{U}^{\oplus n}$ for which the collection $(\mathcal{O}_{U}^{\oplus n},\bigoplus_{j=1}^{n}\nabla_{d^{[j]}},\mathfrak{f}^{\prime}_{\circ})$ forms a parabolic $p^{N}$-flat bundle. The quasi-parabolic structure $\mathfrak{f}_{\circ}$ on $\mathcal{F}$ corresponding to $\mathfrak{f}^{\prime}_{\circ}$ via $\xi$ provides a required one.
Moreover, since $a_{\circ}^{[1]}<\cdots<a_{\circ}^{[n]}$, it follows from the definition of $\mathfrak{f}^{\prime}_{\circ}$ and  [Wak4, Proposition 4.3.4, (ii)] that $\mathfrak{f}^{\prime}_{\circ}$ is invariant under transposing by any automorphism of $\bigoplus_{i=1}^{n}\mathscr{O}_{d_{\circ}^{[j]},\flat}$. Hence, $\mathfrak{f}_{\circ}$ does not depend on the choice of $\xi$. This proves the uniqueness assertion, as desired. ∎
@@
Let us keep the above notation and assumptions, and suppose that $\sum\limits_{j=1}^{n}a_{i}^{[j]}<p^{N}$ for every $i=1,\cdots,r$. Then, the following assertions hold.
@@
Let us take a line bundle $\mathcal{L}$ on $X^{(1)}$ of even degree and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in(\Xi_{2,1}^{<})^{\times r}$, where $a_{i}:=(a_{i}^{[1]},a_{i}^{[2]})$ ($i=1,\cdots,r$). We write $\omega:=(2,\vec{1}_{\times 2},\vec{a}/p)$, and suppose that $r+\sum\limits_{i=1}^{r}(a_{i}^{[1]}+a_{i}^{[2]})$ is even and $\sum\limits_{i=1}^{r}(a_{i}^{[2]}-a_{i}^{[1]})<2g-2+r<\frac{p}{2}$. Then, the cardinality $\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))$ of the set of isomorphism classes of objects in $\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k)$ satisfies the following inequality:
@@
where, for an integer $b$ with $0\leq b\leq p-1$, we set $\tau(b):=\frac{b-1}{2}$ (resp., $\tau(b):=\frac{p-1-b}{2}$) if $b$ is odd (resp., even). On the other hand, the following equality holds: